OpenStudy (anonymous):

Determine where the function is concave upward and downward, and list all the - value(s) of the point(s) of inflection: f(x)=4x^3+2x^2+2x-9

OpenStudy (anonymous):

where are you stuck on this one?

OpenStudy (anonymous):

i got (-inf,-1/6) (-1/6,inf) whats the inflections point

OpenStudy (anonymous):

for the inflection point you plug your critical point from the second derivative back into the second derivative equation - since the inflection point is where the concavity changes.

OpenStudy (anonymous):

thus f'(x) = 12x^2+4x+2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so f'' is just 24x+4 ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

set the second derivative to zero to get the critical point ?

OpenStudy (anonymous):

-1/6

OpenStudy (anonymous):

plug that back into f(x) sorry i said second deriv earlier. but if you want coordinates.. you just have to plug back into f(x)

OpenStudy (anonymous):

i did but for some reason my answer is coming up wrong

OpenStudy (anonymous):

sec working it on paper myself. sorry.

OpenStudy (anonymous):

nevermind i got it! it was only asking for the x values i was putting in the point

OpenStudy (anonymous):

do you mind helping me with another problem

OpenStudy (anonymous):

ooh lol.

OpenStudy (anonymous):

not at all.. go ahead.

OpenStudy (anonymous):

Suppose that the position of a particle as a function of time (in seconds)is given by the formula s(t)=5.5+5t^3-t^5 t>0 find The time at which the velocity is maximum

OpenStudy (anonymous):

ok so same thing here. the the derivative of s(t) to get velocity right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then find the critical points of the derivitive

OpenStudy (anonymous):

i got x= -sqrt(3) x=0 and x=sqrt(3)

OpenStudy (anonymous):

yes that is the velocity. since the domain is x>0 the negative number is not what we want.

OpenStudy (anonymous):

i already fouund velocity and acceleration.. im having trouble finding when velocity is maximum

OpenStudy (anonymous):

when you find the critical point for the first derivatives.. its either a local max, local min, abs max, or abs min.

OpenStudy (anonymous):

so the max is going to occur when v(t)=0 or t=sqrt(3) as you said.

OpenStudy (anonymous):

v(0)=5.5 and v(sqrt(3) is 15.8923

OpenStudy (anonymous):

you only need to set the velocity v(t) = 0. which you did at got 1.73.... or sqrt3. so @ time 1.73 seconds, the particle is moving at its maximum velocity. if you need the position then you plug sqrt3 back into s(t).

OpenStudy (anonymous):

its coming up wrong

OpenStudy (anonymous):

you using MML?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry if im not much help.. im also a calc student this semester. Do you think if we set the acceleration = to 0 we can see what time the velocity is changing? thus my thinking is that when a(t)=0, the velocity is at a max? maybe it isn't right. i got t=1.22 seconds. I suppose it would help to graph it. I have not done that.

OpenStudy (anonymous):

its alright, ill just skip that part for now

OpenStudy (anonymous):

@Luis_Rivera might be able to help.. he is good.

OpenStudy (anonymous):

thanks